How do you solve:
A painter can paint a certain house in 12 days. Working with his partner, the two can paint the house in 4 days. How long would it take the partner working alone to paint the house?
Thank you!

Found 2 solutions by nerdybill, vksarvepalli:
Answer by nerdybill(7384)

A painter can paint a certain house in 12 days. Working with his partner, the two can paint the house in 4 days. How long would it take the partner working alone to paint the house?
.
Let x = number of days for partner to paint house by himself
then
4(1/12 + 1/x) = 1
Multiplying both sides by 12x:
4(x + 12) = 12x
4x + 48 = 12x
48 = 8x
6 days = x

Answer by vksarvepalli(154)

A painter can paint a certain house in 12 days
so in one day he can paint 1/12 th of the house
Working with his partner, the two can paint the house in 4 days
so in one day they both together can paint 1/4th of the house
let the partner paint 1/x th of the house in one day.

so 1/12 + 1/x = 1/4
=> 1/x = 1/4 - 1/12
=> 1/x = 1/6
therefore x = 6
so the partner can paint 1/6 th of the house in one day

so he will take 6 days to paint the house completely